OpenStudy (buttermequeasy):

Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (anonymous):

Divide by -5 |x + 1| = -2 x + 1 = -2 => x = -3 x + 1 = 2 => x = 1